chuckthedonut:

This is the last question I had trouble with- As much help is appreciated. Thank you <3

chuckthedonut:

@supie and if you can help again (last time) @oliver69 that would be nice <3

OLIVER69:

(First Question) So in order to find the exact area you need the equation first which is A = π(11)^2 Then solve it step by step A = π(121) A = 380.13

OLIVER69:

(Second Question) You do the same as the first one but plug 3.14 in for pi. A = 3.14(121) A = 379.94

chuckthedonut:

@oliver69 wrote:
(Second Question) You do the same as the first one but plug 3.14 in for pi. A = 3.14(121) A = 379.94
Ohhhh, I'm surprised that I understand this more than I did in the beginning, tysm! <3

chuckthedonut:

@chuckthedonut wrote:
@oliver69 wrote:
(Second Question) You do the same as the first one but plug 3.14 in for pi. A = 3.14(121) A = 379.94
Ohhhh, I'm surprised that I understand this more than I did in the beginning, tysm! <3
It makes sense how the answers are that, tysm!

OLIVER69:

you're welcome and I'm glad to help. < 3